OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 2x2 + 4x + 3 = 0.

OpenStudy (anonymous):

@myko Can you help?

OpenStudy (anonymous):

discriminant of 2º order equation: \[ax ^{2} +bx +c =0\] is:\[D=b ^{2}-4ac\] depending on the sign of D you know what are the roots: D>0 : 2 distinct real roots D=0: 2 equal real roots D<0: no real roots

OpenStudy (anonymous):

in your case a=2, b=4, c=3

OpenStudy (anonymous):

So it's 2 equal real roots?

OpenStudy (anonymous):

That's what I'm getting @myko

OpenStudy (anonymous):

D=16-4*2*3=16-24=-8 no real roots

OpenStudy (anonymous):

Oh, I see what I did wrong.